I am a7 grade student and I do not understand how to work these problems can someone do astep by step for thank yuo very much
Find the missing number,n
185,86,79,140,192,36,185,85,85,112,n mean=115
129,60,63,52,27,n,75,193,67 mean=77
please help I need this tomorrow.

Answer by checkley71(8403)

185+86+79+140+192+36+185+85+85+112+N=(MEAN OF 115)
YOU ARE LOOKING FOR THE AVERAGE OF ALL 11 NUMBERS THUS YOU ADD THEM ALL UP & DIVIDE BY 11.
(1185+X)/11=115 NOW WE HAVE TO CROSS MULTIPLY TO GET RID OF THE DENOMINATOR
1185+X=1265 NOW WE SUBTRACT 1185 FROM BOTH SIDES
X=1265-1185
X=80 IS THE MISSING NUMBERPROOF
(1185+80)/11=115
1265/11=115
115=115